OpenStudy (anonymous):

If f(x) = │(x² – 50)│, what is the value of f(-5) ? A. 75 B. 25 C. 0 D. -25 E. -75 Walk me through it please. Don't just give me the answer

OpenStudy (anonymous):

They said F(x) later and they give me F(-5) so would i replace -5 with x?

OpenStudy (igreen):

\(f(x) = |(x^2 - 50)|\) \(f(-5) = |(-5^2 - 50)|\) I think you can finish that.

OpenStudy (igreen):

Yep, you do.

OpenStudy (anonymous):

D. -25

OpenStudy (anonymous):

also i meant to say replace x with -5 :p

OpenStudy (igreen):

Close. It has absolute value signs..which makes the value positive. \(f(-5) = |(-5^2 - 50)|\) \(f(-5) = |(25 - 50)|\) \(f(-5) = |-25|\)

OpenStudy (anonymous):

Oh okay so B. 25 thank you

OpenStudy (igreen):

Yep.